OpenStudy (anonymous):

Write a radical in which the radicand consists of a coefficient and a variable. Use a coefficient that can be simplified best using the greatest perfect square method. Then, simplify the radical completely. You can use "sqrt()" to show a square root. For example, can be written as 7sqrt(a).

OpenStudy (anonymous):

alright so you want to basically pick a perfect square coefficient

OpenStudy (anonymous):

these are \[1^2,2^2,3^2\] ... the reason why you want to do this is because a perfect square can be taen out of a radical sign . For example \[\sqrt{9}=\sqrt{3*3} \][if you have 2 within the square root, you can take the three out \[\sqrt{9}=3\]

OpenStudy (anonymous):

Alright so now pick a perfect square coefficient(number) and set it within the radical with a variable

OpenStudy (anonymous):

could you solve it for me please

OpenStudy (anonymous):

@Outkast3r09